### v1

@@ 4.1 Datasets
The DVM dataset comprises 35,719 image-tabular pairs, each consisting of a 300$\times$300 front-view RGB image of a car and corresponding tabular data on its sales and technical data. To avoid label leakage in the car model classification task, we excluded features such as brand, model year, and vehicle size, resulting in 13 tabular input features. Following the protocol in (Hager et al., 2023), we considered 101 target classes, each with at least 100 samples. Top-1 accuracy (ACC) was used as the evaluation metric, and the data was randomly split into training, validation, and test sets with a ratio of 72%, 18%, and 10%, respectively. Additional preprocessing details are provided in Appendix A.1.
4.2 Experimental Settings
4.2.1 Evaluating Tabular Encoders in Supervised Learning
We first compared four tabular encoders, FT-Mamba, MLP (Hager et al., 2023), ResNet (Gorishniy et al., 2021), and FT-Transformer (Gorishniy et al., 2021), by training unimodal tabular models under supervised learning. Each model consisted of the corresponding encoder followed by an FC classification layer. The number of encoder blocks was fixed to 2 across all models, following the configurations in (Hager et al., 2023; Gorishniy et al., 2021). Based on hyperparameter search, the tabular embedding dimension $D_{tab}$ was set to 256 for MLP and ResNet, and 64 for FT-Transformer and FT-Mamba. Results are reported as the mean and standard deviation across five different random seeds. Additional implementation details are provided in Appendix A.3.
To investigate how the tabular encoders process high-dimensional tabular inputs, we performed Integrated Gradient (IG) analysis on the OL3I dataset. IG estimates the contribution of each input feature by integrating gradient values along a straight-line path from a baseline input to the target sample (Sundararajan et al., 2017). In our setting, the baseline input was defined as the mean of all training tabular samples. IG scores were computed for all training samples with respect to the disease class, and the final feature importance values were averaged to mitigate randomness. For consistent comparison, each feature’s importance was normalized as a percentage of the total importance within the model, thereby highlighting its relative influence during inference.
4.2.2 Multimodal Contrastive Pretraining and Interpretation
We selected ResNet-50 as the image encoder and pretrained it in combination with each of the four tabular encoders introduced in Section 4.2.1, resulting in four distinct image-tabular encoder pairs. All models were trained from scratch using contrastive learning. An exception was made for the OL3I experiment involving ResNet-50 and FT-Transformer, where the contrastive loss failed to converge; in this case, we initialized the ResNet-50 with weights from the ResNet-50-MLP pretraining. Additional implementation details are provided in Appendix A.2.
To investigate how different tabular encoders influence the image modality under contrastive learning, we applied Grad-CAM (Selvaraju et al., 2017) to generate class activation maps (CAMs) from the pretrained multimodal models. For a consistent comparison, all models used a concatenation-based fusion framework to avoid confounding effects from fusion modules.
4.2.3 Multimodal Fine-tuning for Downstream Tasks
In the fine-tuning stage, we compared three multimodal fusion methods: concatenation, cross-attention, and the proposed AMF module. In the concatenation fusion method, we followed the approach in (Hager et al., 2023), where the unimodal feature vectors generated by the pretrained encoders were concatenated and fed into an FC layer for classification. In the cross-attention fusion method, we followed the approach in (Du et al., 2024). The feature map from the final layer of ResNet-50 was flattened into image tokens. These tokens, along with the tabular token generated by FT-Transformer or FT-Mamba, were then fed into a multimodal cross-attention module with 2 layers and an embedding dimension of 512. Finally, the CLS token was passed to the classification layer for prediction. In the AMF fusion, the configurations of modality confidence ratio $R_{conf}$ are provided in Appendix A.4.
We applied both frozen and trainable strategies to each pretrained model during fine-tuning. In the frozen strategy, all parameters of the pretrained encoders were frozen. In the retraining strategy, all weights in the model were trainable. A unimodal ResNet-50 was also trained with supervised learning for comparison. More parameter settings can be found in Appendix A.3.
4.2.4 Experiments on Incomplete Tabular Data
Incomplete tabular data is a common challenge in real-world clinical applications. This variability in downstream settings necessitates that robust multimodal models remain effective even when one modality is partially degraded. To evaluate model adaptability under such conditions, we simulated an incomplete tabular input scenario on the OL3I dataset by masking the most important tabular feature, which was replaced with its mean value across all samples. Feature importance was determined using IG analysis based on the supervised tabular models.
Under this incomplete setting, we fine-tuned the multimodal models using a ResNet-50 image encoder and an FT-Mamba tabular encoder, both pretrained via contrastive learning. Three fusion strategies—feature concatenation, cross-attention, and the proposed AMF module with different $R_{conf}$ values—were compared to assess their robustness to missing tabular information.
4.2.5 Ablation Study on the AMF Module
To validate the necessity of dynamically adjusting unimodal feature dimensions based on modality confidence, we conducted an ablation study on the modality confidence ratio $R_{conf}$ in the AMF module. We tested $R_{conf}$ values of 1.0, 0.97, 0.95, and 0.1 during the fine-tuning on the OL3I dataset, which directly controls the allocation of feature dimensions between the image and tabular modality prior to concatenation. Among these, 1.0 and 0.97 correspond to the $R_{conf}$ values selected for ResNet and MLP tabular encoders, respectively, while 0.1 was chosen to simulate a scenario where the dimensionality of tabular features is significantly higher than that of the image features. This setting contrasts with the concatenation-based fusion, where $R_{conf}$ is not applied.
To evaluate the contribution of the proposed auxiliary loss functions—density loss and leakage loss, we fixed $R_{conf}=0.95$ and compared four different loss configurations during fine-tuning: (1) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}$; (2) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}$; (3) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{2}\mathcal{L}_{\text{density}}$; (4) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}$ (ours).
4.3 Unimodal Performance of Tabular Models
4.3.1 Performance
Table 1: Performance of unimodal tabular models trained with supervised learning.
Encoder OL3I AUC (%) DVM ACC (%)
MLP 74.67 $\pm$ 2.06 85.50 $\pm$ 0.50
ResNet 74.43 $\pm$ 1.08 85.19 $\pm$ 0.37
FT-Transformer 74.66 $\pm$ 1.25 94.30 $\pm$ 0.29
FT-Mamba 77.36 $\pm$ 1.58 92.71 $\pm$ 0.28
The unimodal performance of the four tabular encoders under supervised training is shown in Table 1, with the best results highlighted in bold. Overall, FT-Mamba demonstrates strong performance on both datasets, particularly excelling on the medical dataset, while FT-Transformer shows a clear advantage only on the large-scale DVM dataset. On the medical OL3I dataset, FT-Mamba significantly outperforms the other encoders, achieving an average AUC of 77.36%. On the DVM dataset, which contains a larger number of samples, FT-Transformer achieves the highest average accuracy of 94.30%, followed closely by FT-Mamba with 92.71%. In contrast, MLP and ResNet perform substantially worse on both datasets.
4.3.2 Integrated Gradient Analysis
@@
Figure 4: Feature importance distribution for unimodal tabular models on the OL3I dataset.
@@
Figure 5: IG analysis results for unimodal tabular models. (a): Feature importance proportions by category for the OL3I dataset. Features with an importance percentage less than 1% were classified as low-importance features (light blue). Features with an importance percentage of 1% or greater were classified as high-importance features. Based on (Zambrano Chaves et al., 2023) and established medical knowledge, we further divided the high-importance features into medically relevant features (pink) and low-relevance features (dark blue). (b) and (c): High-importance features and their importance proportion for FT-Mamba and FT-Transformer. (exp_wt_lab_systolic: exponentially weighted average of systolic blood pressure. CPT: Current Procedural Terminology codes, ATC: Anatomical Therapeutic Chemical codes.)
Our IG analysis on the OL3I dataset further interprets the behavior of the four tabular encoders. Fig. 4 shows the overall distribution of feature importance across all tabular features. Due to the high dimensionality of the tabular data, we additionally grouped the 423 features into three categories and visualized their respective importance distributions in Fig. 5(a). More detailed IG results are available in Appendix B.
The results indicate that MLP and ResNet exhibit limited capacity for distinguishing tabular features. As illustrated in Fig. 4(a), MLP tends to distribute its attention relatively uniformly across all input features. Although ResNet demonstrates a single peak in its importance distribution, Table B.1 shows that its most important features account for only 18% of the total importance—substantially lower than the 45% observed in FT-Mamba and the 36% in FT-Transformer. Furthermore, as shown in Fig. 5(a), more than 60% of ResNet and MLP’s attention is dispersed across a large number of low-importance features.
In contrast, FT-Mamba and FT-Transformer are more effective at identifying meaningful features from high-dimensional tabular data. Over 60% of their attention is highly concentrated on a small subset of high-importance features, while other features receive minimal attention. This strong feature discrimination capacity contributes to their superior performance when processing high-dimensional inputs.
However, further comparing the FT-Mamba and FT-Transformer, the high-importance but low-relevance-in-medicine features of FT-Transformer account for 10.63%, which is higher than FT-Mamba and even ResNet. The specific high-importance features for them are presented in Fig. 5(b) and 5(c). All of the high-importance features for FT-Mamba are strongly medically relevant to IHD, including cardiovascular risk factors such as age, CCI, and sex, which have been validated as the most important features for IHD prediction in (Zambrano Chaves et al., 2023), along with other records related to the cardiovascular system. In contrast, FT-Transformer places significant attention on several irrelevant features, such as A12, A06, and J05. This erroneous focus shows that FT-Transformer captures several noise patterns when the training data is limited, which may be the factor for the inferior performance of FT-Transformer compared to FT-Mamba on the OL3I dataset.
4.4 Multimodal Model Evaluation on Complete Tabular Data
4.4.1 Model Interpretation in Contrastive Learning
Fig. 6 shows the CAMs generated by four pretrained image-tabular encoder pairs, along with a unimodal ResNet-50 trained using supervised learning. The CAMs generated by the supervised ResNet-50 lack distinct focal regions, indicating insufficient supervision and limited ability of the image encoder to capture meaningful visual features. CAMs from models using MLP or ResNet as tabular encoders show concentrated attention around the lumbar spine, exhibiting a localized and singular focused pattern. In contrast, models with FT-Mamba and FT-Transformer encoders show a more flexible, multi-point attention distribution.
@@ 4.4.1 Model Interpretation in Contrastive Learning
These findings are consistent with the tabular IG analysis in Section 4.3.2 and reflect how different tabular encoders guide the image modality in contrastive learning. ResNet-50, as a convolutional network, inherently emphasizes local feature extraction and is limited in global modeling. When paired with relatively weak tabular encoders like MLP and ResNet, the supervision signal is limited, resulting in minimal changes in the spatial focus of the image encoder. In contrast, FT-Transformer and FT-Mamba provide stronger global supervision through their advanced long-range modeling capacities, enabling the image encoder to learn more structured and informative visual representations.
However, the model with FT-Transformer exhibits suboptimal attention, with substantial activation over irrelevant or blank regions, suggesting that its noisier feature patterns misdirected the image encoder during contrastive training. In contrast, the model with FT-Mamba effectively attends to clinically relevant anatomical structures, including the surrounding musculature, subcutaneous and visceral fat, and cardiovascular regions. Notably, visceral fat proportion is a key biomarker in predicting IHD (Zambrano Chaves et al., 2023). These results highlight FT-Mamba’s superior capability to support informative feature learning in the image modality, especially under the limited data conditions of the OL3I dataset.
4.4.2 Multimodal Performance on Downstream Classification
The performance of the multimodal models on the test set is summarized in Table 2. All models utilizing trainable fusion strategies substantially outperform the unimodal ResNet-50 trained solely on image data. Among all combinations of fusion methods and tabular encoders, our proposed model—AMF module combined with FT-Mamba—achieves the highest AUC scores on the OL3I dataset under both frozen and trainable strategies (76.97% and 78.36%, respectively). On the large-scale DVM dataset, the best results are obtained by the cross-attention + FT-Mamba model in the frozen setting (96.96% ACC), and by the AMF + FT-Transformer model in the trainable setting (99.23% ACC).
Table 2: Classification performance of multimodal methods on complete tabular data.
Method OL3I AUC (%) DVM ACC (%)

### v2

@@ 4.1 Datasets
The DVM dataset comprises 35,719 image-tabular pairs, each consisting of a $\times300$ front-view RGB image of a car and corresponding tabular data on its sales and technical data. To avoid label leakage in the car model classification task, we excluded features such as brand, model year, and vehicle size, resulting in 13 tabular input features. Following the protocol in [6], we considered 101 target classes, each with at least 100 samples. Top-1 accuracy (ACC) was used as the evaluation metric, and the data was randomly split into training, validation, and test sets with a ratio of 72%, 18%, and 10%, respectively. Additional preprocessing details are provided in Appendix A.1.
4.2 Experimental Setup
In our experiments, the image encoder is a ResNet-50, which generates image features of dimension 2048. The tabular encoder is implemented as FT-Mamba, consisting of two Mamba blocks, and produces tabular features of dimension 64. All projection heads in the network are MLPs with a single hidden layer. During pretraining, the projection heads map features to a 128-dimensional space [6]. In the AMF module, the projection heads project features to a 2048-dimensional space. A fully connected layer is used for the final classification.
We first pretrain the image and tabular encoders on the OL3I and DVM datasets, respectively, following the contrastive learning strategy described in [6]. Prior to fine-tuning, each pretrained encoder is frozen and paired with an FC classification layer to form a unimodal model. The unimodal performance on the downstream task is then used to compute the modality confidence ratio $r_{\text{conf}}$ according to (2). For downstream tasks, fine-tuning is performed under two strategies: in the frozen strategy, the parameters of pretrained encoders remain fixed, whereas in the trainable strategy, all parameters are updated during training. All results are reported as the mean and standard deviation across five different random seeds, and more implementation details are provided in Appendix A.
4.3 Comparison with Recent Methods on Downstream Classification
To evaluate the effectiveness of our proposed AMF-MedIT framework, we compare it with both supervised baselines and recent SSL approaches on the two downstream classification tasks. The supervised baselines include a ResNet-50 for the image modality [28] , an FT-Transformer for the tabular modality [25], and a Concat Fuse method for medical image and tabular fusion [33]. For SSL-based multimodal methods, we include MMCL [6], TIP [13], and AMF-MedIT. MMCL and AMF-MedIT follow the same pretraining strategy, while TIP adopts the pretraining procedure described in its original paper.
Table 1: Classification performance of AMF-MedIT and recent methods for image and tabular data.
Model OL3I AUC (%) DVM ACC (%)
[…]
Supervised Unimodal and Multimodal Methods
ResNet-50 [28] 58.29 $\pm$ 4.16 93.00 $\pm$ 1.05
FT-Transformer [25] 74.66 $\pm$ 1.25 94.30 $\pm$ 0.29
Concat Fuse [33] 76.49 $\pm$ 0.61 90.37 $\pm$ 7.54
SSL Multimodal Pre-training Methods
MMCL [6] 76.31 $\pm$ 0.45 77.18 $\pm$ 0.82 93.51 $\pm$ 1.06 97.45 $\pm$ 0.19
TIP [13] 73.02 $\pm$ 0.40 75.87 $\pm$ 0.44 98.84 $\pm$ 0.05 97.91 $\pm$ 0.11
AMF-MedIT (ours) 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11
  • 1
    Frozen strategy.
  • 2
    Trainable strategy.
The results are summarized in Table 1. For the OL3I task, AMF-MedIT achieves the best performance under the trainable strategy, with an AUC of 78.36%, outperforming both MMCL (77.18%) and TIP (75.87%). For the DVM task, which contains a substantially larger number of training samples, AMF-MedIT also achieves the highest accuracy of 98.89%, marginally higher than TIP (98.84%). While TIP can fully exploit its attention-intensive architecture and achieve competitive performance when ample data are available, its reliance on large datasets limits its generalization to smaller medical cohorts. In contrast, AMF-MedIT shows consistent advantages across both large- and small-scale settings, benefiting from its straightforward yet adaptive fusion design. Furthermore, compared with supervised unimodal baselines, the improvements are substantial, underscoring the advantages of multimodal SSL pretraining.
4.4 Robustness to Noisy Data
To evaluate the robustness of the models under noisy conditions, we simulated two types of noise commonly encountered in clinical practice based on the OL3I dataset: tabular data missingness and additive Gaussian noise in images.
  • Tabular data missingness: We adopted the random feature missingness (RFM) strategy, where a random subset of feature columns is removed and the missing values are imputed with the mean of the corresponding feature. The missing rates were set to $[0,0.25,0.5,0.75]$.
  • Image Gaussian noise: We injected additive Gaussian noise into the images, with noise levels controlled by the standard deviation $\sigma$, set to $[0,0.1,0.15,0.25]$.
We designed two experimental scenarios: (i) noisy tabular data with clean images, and (ii) noisy images with clean tabular data. In each scenario, the noise was applied to the training, validation, and test splits of the downstream tasks, and we fine-tuned the pretrained models of three image-tabular SSL frameworks (MMCL [6], TIP [13], and ours) on the noisy datasets and compared their downstream performance. Fig. 4 presents the multimodal performance on noisy data.
(a) Performance on missing tabular data and clean image data
(b) Performance on noisy image data and clean tabular data
Figure 4: Performance comparison with SSL image-tabular methods on noisy data.
For tabular missingness, TIP achieves relatively low absolute AUC but shows the smallest performance drop, indicating strong stability. This advantage comes from its dedicated missing-feature processing mechanism, which masks absent tabular features and enables consistent adaptation under both frozen and trainable settings. Although AMF-MedIT demonstrates strong performance and good adaptability under the frozen setting, its performance degrades notably under the trainable setting when the missing rate exceeds 0.5. This limitation may arise because $r_{\text{conf}}$ is configured based on the frozen-encoder performance in our experiments; at high missing rates, the confidence scores of frozen and trainable encoders diverge, leading to suboptimal adjustment. In future work, the settings of $r_{\text{conf}}$ could be refined by leveraging unimodal performance under the trainable strategy to achieve a more accurate estimation, and even further adjusted by incorporating domain expertise from medical professionals.
For image noise, AMF-MedIT delivers the best overall performance with minimal degradation, demonstrating strong adaptability. TIP, however, is less effective in this case. Since its mechanism is tailored only for handling tabular missingness, it cannot accommodate degraded image quality, resulting in weaker robustness compared with AMF-MedIT and MMCL.
Overall, AMF-MedIT exhibits the strongest robustness across both noise scenarios. Rather than depending on a modality-specific mechanism, the $r_{\text{conf}}$ adjustment in our model operates in a modality-agnostic manner, providing broader adaptability.
4.5 Ablation Studies
4.5.1 Ablation Studies on Tabular Encoder and Model Interpretation
To evaluate the effectiveness of the proposed FT-Mamba tabular encoder within AMF-MedIT, we replace it with three representative tabular encoders: MLP [6, 33], ResNet [25], and FT-Transformer [25]. Each encoder is trained under three settings: (i) unimodal supervised training, (ii) multimodal contrastive pretraining, and (iii) multimodal fine-tuning on downstream tasks. To further interpret how different tabular encoders operate and influence contrastive learning, we additionally perform Integrated Gradients (IG) analysis [31] on unimodal tabular models and Grad-CAM analysis [32] on multimodal encoders.
Performance.
The results of tabular encoders on the OL3I and DVM datasets are summarized in Table 2. On the OL3I dataset, FT-Mamba achieves the best AUC both in unimodal (77.36%) and multimodal (78.36%) settings. On the DVM dataset, FT-Transformer achieves the best overall accuracy in both unimodal supervised training (94.30%) and multimodal fine-tuning (99.23%), with FT-Mamba ranking closely behind (92.71% and 98.89%, respectively). In contrast, MLP and ResNet performed much worse. Overall, these results suggest that FT-Mamba offers better adaptability across datasets of different scales: it shows clear advantages on the smaller OL3I dataset, while remaining competitive with FT-Transformer on the larger DVM dataset.
Table 2: Ablation studies on different tabular encoders.
Tabular Encoder OL3I AUC (%) DVM ACC (%)
[…]
MLP [6] 74.67 $\pm$ 2.06 76.63 $\pm$ 0.22 76.86 $\pm$ 0.35 85.50 $\pm$ 0.50 94.92 $\pm$ 0.43 97.88 $\pm$ 0.42
ResNet [25] 74.43 $\pm$ 1.08 76.50 $\pm$ 0.83 75.46 $\pm$ 0.92 85.19 $\pm$ 0.37 95.09 $\pm$ 0.54 97.28 $\pm$ 0.70
FT-Transformer [25] 74.66 $\pm$ 1.25 75.62 $\pm$ 0.49 76.46 $\pm$ 1.57 94.30 $\pm$ 0.29 93.83 $\pm$ 0.90 99.23 $\pm$ 0.35
FT-Mamba (ours) 77.36 $\pm$ 1.58 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 92.71 $\pm$ 0.28 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11
Integrated gradients analysis of tabular encoders.
Fig. 5 shows the overall distribution of feature importance estimated by IG across all tabular features on the OL3I dataset, while Fig. 6(a) groups the 423 features into three clinical categories for clarity. More detailed IG results are provided in Appendix B. The results reveal that MLP and ResNet show weak feature discrimination: MLP distributes attention almost uniformly across features, while ResNet exhibits a single peak but still spreads over many irrelevant variables. Quantitatively, their most important features account for only 18% (ResNet) and even less for MLP of the total importance, far below FT-Mamba (45%) and FT-Transformer (36%) according to Table B.1. As a result, over 60% of their attention remains dispersed across low-importance features, reflecting a limited ability to extract clinically meaningful signals.
@@ 4.5.1 Ablation Studies on Tabular Encoder and Model Interpretation
Figure 5: Importance distribution over 423 input features for unimodal tabular models.
@@
Figure 6: IG analysis results for unimodal tabular models. (a): Feature importance proportions by category for the OL3I dataset. Features with an importance percentage less than 1% were classified as low-importance features (light blue). Features with an importance percentage of 1% or greater were classified as high-importance features. Based on [4] and established medical knowledge, we further divided the high-importance features into medically relevant features (pink) and low-relevance features (dark blue). (b) and (c): High-importance features and their importance proportion for FT-Mamba and FT-Transformer. (exp_wt_lab_systolic: exponentially weighted average of systolic blood pressure. CPT: Current Procedural Terminology codes, ATC: Anatomical Therapeutic Chemical codes.)
In contrast, FT-Mamba and FT-Transformer concentrate on a small subset of features, with more than 60% of their attention massed on clinically salient variables. However, the patterns differ in quality. Fig. 6(b) and 6(c) show the specific high-importance features for FT-Mamba and FT-Transformer. FT-Mamba’s high-importance features are strongly aligned with well-established cardiovascular risk factors for IHD, such as age, sex, Charlson Comorbidity Index (CCI), and records related to the cardiovascular system, which are consistent with the prior findings [4]. By comparison, FT-Transformer assigns substantial attention to features with little medical relevance (e.g., A12, A06, J05), suggesting it may capture noise patterns when training data are limited.
Grad-CAM analysis of multimodal encoders.
Fig. 7 visualizes Grad-CAMs from different image–tabular encoder pairs, alongside a unimodal supervised ResNet-50. The supervised ResNet-50 shows no distinct focal regions, confirming its limited capacity to capture meaningful visual cues under scarce supervision. Models with MLP or ResNet tabular encoders produce narrow attention localized around the lumbar spine, reflecting weak and overly singular guidance to the image encoder. In contrast, FT-Mamba and FT-Transformer enable broader, multi-point attention patterns, consistent with their stronger long-range modeling ability. However, FT-Transformer often activates irrelevant or blank areas, suggesting its noisier tabular features misguide the image encoder during contrastive training. By comparison, FT-Mamba directs attention toward clinically relevant structures—including musculature, subcutaneous and visceral fat, and cardiovascular regions—among which visceral fat proportion is a key biomarker in predicting IHD [4].
@@
Taken together, these results highlight the flexible and efficient capability of FT-Mamba in modeling tabular data. It can effectively extract highly relevant tabular features without being constrained by limited data or abundant irrelevant variables. Furthermore, it demonstrates superior capacity to provide medically meaningful supervision for the image modality. This simple yet effective working mechanism enables AMF-MedIT to achieve advantages in learning both tabular and image feature representations.
4.5.2 Ablation Study on Multimodal Fusion Module
To evaluate the effectiveness of the proposed AMF module, we replaced it with two mainstream multimodal fusion methods: concatenation and cross-attention. In the concatenation approach [6, 33], the unimodal feature vectors generated by the pretrained encoders are concatenated and fed into a fully connected layer for classification. For cross-attention [13], the feature map from the final ResNet-50 layer is flattened into image tokens, which are then combined with the tabular token generated by FT-Transformer or FT-Mamba and passed through a 2-layer multimodal cross-attention module with an embedding dimension of 512. The resulting CLS token is finally used for prediction. Table 3 summarizes both classification performance and computational costs. All measurements were conducted with a batch size of 1. The inference latency was measured on an NVIDIA RTX 4090 GPU, averaged over 100 runs.
Table 3: Ablation studies on different multimodal fusion modules.
Fusion Module #Param (M) OL3I AUC (%) OL3I (Cost) DVM ACC (%) DVM (Cost)